OpenStudy (anonymous):

simplify:

OpenStudy (anonymous):

\[5\sqrt{3}(2\sqrt{3}+6\sqrt{19})\]

OpenStudy (primeralph):

expand

OpenStudy (anonymous):

do you any answer choices?

OpenStudy (anonymous):

yea hold on

OpenStudy (anonymous):

\[30+30\sqrt{57}\] \[90+30\sqrt{57}\] \[10\sqrt{3}+30\sqrt{19}\] \[10\sqrt{3}+30\]

OpenStudy (anonymous):

@KeganMoore

OpenStudy (anonymous):

Go with @allie_bear22 lol

OpenStudy (anonymous):

i didnt write anything lol

OpenStudy (anonymous):

haha bro my girl need help here lol @KeganMoore

OpenStudy (anonymous):

then what is with the equations you typed out?

OpenStudy (anonymous):

those were the choices

OpenStudy (anonymous):

5√3(2√3+6√19) @KeganMoore

OpenStudy (anonymous):

@Hope_nicole could use some help doing this for them

OpenStudy (anonymous):

\[5 \sqrt{3} \left(2 \sqrt{3}+6 \sqrt{19}\right)=30+30 \sqrt{57} \]

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

@robtobey and how did you get that? its easy to post an answer but explanations are better

OpenStudy (anonymous):

yup mr robeert!

OpenStudy (anonymous):

tru

OpenStudy (anonymous):

he distributed the \[5\sqrt{3}\]to the rest of the equation.

OpenStudy (anonymous):

if you don't explain a question how do you expect them to understand it?

OpenStudy (anonymous):

when he did that, the \[5\sqrt{3} and 2\sqrt{3} \because 5*2*3 since the two \sqrt{3} would become 3\]

OpenStudy (anonymous):

@KeganMoore \[a (b+c)=a b+a c \]\[5\sqrt{3}\left(2\sqrt{3}\right)=10*3 \]\[5\sqrt{3}6\sqrt{ 19}=30\sqrt{3}\sqrt{ 19}=30\sqrt{3*19} \]

OpenStudy (anonymous):

ok sweet...i have a learning disorder primarily in math

OpenStudy (anonymous):

Mistake in the 3rd expression. Sould be:\[5\sqrt{3}\left(6\sqrt{ 19}\right)=30\sqrt{3}\sqrt{ 19}=30\sqrt{3*19} \]Sorry about that.